OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (mathstudent55):

Let b = number of boys Let g = number of girls There are 89 boys and girls combined, so b + g = 89 "the number of boys is 8 more than twice the number of girls" b = 8 + 2 * g b = 2g + 8 The two equations are b + g = 89 b = 2g + 8 Now you need to solve the system of equations.

OpenStudy (anonymous):

Its D.

OpenStudy (mathstudent55):

b + g = 89 b = 2g + 8 Solve the first equation for g and substitute into second equation: g = 89 - b b = 2(89 - b) + 8 b = 178 - 2b + 8 3b = 186 b = 62 You are correct.